OpenStudy (anonymous):

What is the value of x in the equation 2 x + 7 = 1?

OpenStudy (cggurumanjunath):

2x+7 =1 2x = 1 - 7 = -6 x =-6/2

OpenStudy (cggurumanjunath):

x = ? @Aliypop

OpenStudy (anonymous):

3

OpenStudy (cggurumanjunath):

-3

OpenStudy (anonymous):

What is 1 1/2 x3

OpenStudy (anonymous):

would I divide fist

OpenStudy (anonymous):

first